OpenStudy (anonymous):

Am I setting this problem up correctly? Solve 3^x-4 = 7^ x+9 (x - 4) log 3 = (x + 9) = log 7

OpenStudy (anonymous):

huh :(

OpenStudy (anonymous):

i got 12...

OpenStudy (anonymous):

x = 4 In (3) + 9 In (7)/In (3) - In (7)

OpenStudy (anonymous):

THANK YOU FOR THE STEPS!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

but i still got 12

OpenStudy (anonymous):

i used the calc magic and let me just tell you this....there was no magic!!!!!!!

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (amistre64):

(x - 4) log 3 = (x + 9) log 7 xL3 - 4L3 = xL7 + 9L7 xL3 - xL7 = 4L3 + 9L7 x(L3 - L7) = 4L3 + 9L7 4L3 + 9L7 x = ---------- L3 - L7

OpenStudy (anonymous):

where did L come from???????????????????

OpenStudy (anonymous):

oh Log:)

OpenStudy (amistre64):

im sure it can be cleaned up a little more, but thats the gist of it .... and yes, L for log

OpenStudy (anonymous):

2.38061585769

OpenStudy (amistre64):

... mine says -25.86 or so

OpenStudy (amistre64):

3^x-4 = 7^ x+9 those are exponents right?

OpenStudy (anonymous):

love u